OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (anonymous):

yes it is Pythagorean triple 24^2+7^=25^2

OpenStudy (hoblos):

@kymber !!! 7^2 + 24^2 = 625 = 25^2 so i go with @garryh

OpenStudy (kymber):

What you want to do is plug the numbers that you know into the Pythagorean theorem equation, a² + b² = c² and see if they check out. The largest number will be c, since c is the hypotenuse which is the largest side. So  :D